# Unconditional verification of quantum computation with classical light

### v1

@@ Abstract
Verification of quantum computation is a task to efficiently check whether an output given from a quantum computer is correct. Existing verification protocols conducted between a quantum computer to be verified and a verifier necessitate quantum communication to unconditionally detect any malicious behavior of the quantum computer solving any promise problem in ${\sf BQP}$. In this paper, we remove the necessity of the communication of qubits by proposing a “physically-classical” verification protocol in which the verifier just sends coherent light to the quantum computer.
@@ I Introduction
Quantum computers are expected to outperform classical computers in several applications, from cryptanalysis BHT97 ; S97 ; HSTX20 to physics K95 ; NPdB21 ; CDBBT24 and chemistry KJLMA08 ; GL22 ; CFGHLMW23 . On the flip side of these advantages, they are susceptible to noises. Therefore, to get benefits from quantum computers, it would be necessary to devise an efficient protocol for checking whether a quantum computer outputs a correct answer. This task is called verification of quantum computation GKK19 ; EHWRMPCK20 ; KR21 . Although someone may think that verification protocols become useless if a sufficient number of qubits for quantum error correction G20 will be realized, this is not the case because it will be still necessary to check whether an adopted quantum error correction scheme faithfully works. Multiple small-scale experiments BFKW13 ; ZZCPXYYYHXCLG20 ; JWQCCLXSZM20 ; FMMD21 have already been demonstrated toward the realization of verifiable quantum information processing.
@@
Verification protocols are evaluated in terms of five properties: (i) whether the soundness is information theoretical or computational, (ii) what type of communication is required for a verifier, (iii) how many number of non-communicating provers, which are quantum computers to be verified, are necessary, (iv) the presence or absence of a trusted third party who surely follows procedures of the protocols, and (v) to which problems the protocols can be applied. Here, the information-theoretical and computational soundnesses mean that any computationally-unbounded and quantum-polynomial-time malicious provers who output incorrect answers can be rejected, respectively, and hence the former is a stronger property than the latter. As for (iii), it would, in general, be hard to guarantee that multiple provers do not communicate each other. Therefore, the ultimate goal is to devise a protocol such that (i) the soundness is information theoretical, i.e., even if the quantum computer outputs an incorrect answer by taking superpolynomial time, it can be properly detected, (ii) the classical communication is sufficient, (iii) & (iv) a single prover is sufficient, and (v) the protocol can be applied to any problem in ${\sf BQP}$, which is a set of promise problems (i.e., problems that can be answered by YES or NO) solvable in quantum polynomial time. However, it is hard to construct such outstanding protocol with the currently known theoretical techniques. Although its impossibility was not shown, which immediately implies ${\sf P}\neq{\sf PSPACE}$, some existing results FHM18 ; MT20 reveal the difficulty of its construction.
@@
This situation leaded to several verification protocols M12 ; RUV13 ; M14 ; GKW15 ; HM15 ; MNS16 ; M16 ; HKSE17 ; KW17 ; FK17 ; FHM18 ; M18 ; DOF18 ; TM18 ; B18 ; FKD18 ; LMNT18 ; MTN18 ; TMMMF19 ; CGJV19 ; G19 ; GV19 ; FKD19 ; LDTAF19 ; ZH19 ; M20 ; ACGH20 ; CCY20 ; LMKO21 ; CLLW22 ; KKLMO22 ; M22 ; Z22 ; GMP23 ; LZH23 each of which has own advantage and disadvantage. They fall into five types of approaches FN1 as summarized in Fig. 1. In terms of practicality, we stick to achieving the information-theoretical soundness and keeping the number of participants other than the verifier one. This is because it would be unclear how to guarantee that the prover’s computation is completed in polynomial time, that the multiple provers do not communicate each other, and that a third party does not cooperate with the prover. In this sense, from Fig. 1, quantum communication is necessary to verify any problem in ${\sf BQP}$ with existing practical verification protocols.
@@
Figure 1: Summary of existing verification protocols. The third column indicates what type of communication is required for the verifier. In the fifth column, the cross and circle marks mean that a trusted third party is not necessary and necessary, respectively. In the sixth column, BQP means that the protocols can be applied to any problem in ${\sf BQP}$. The disadvantages of the verification protocols are indicated by bold red words. Our contribution is emphasized by the bold blue word.
@@
In this paper, we remove the necessity of the communication of qubits by proposing a “physically-classical” verification protocol in which the verifier just sends coherent light to the quantum computer to be verified. The transmission of coherent light is the same as classical communication in the sense that classical bits are sent by using light in the real world, while they are absolutely different from the viewpoint of information theory. To obtain our result, we first modify the verification protocol in Ref. M20 such that a trusted third party is combined with the verifier, and hence communication of qubits becomes necessary for the verifier. Then, to remove the communication of qubits, we combine it with a technique of using coherent light as a substitute for qubits. So far, such technique has been developed in several quantum information processing tasks such as quantum key distribution (QKD) B92 ; HIGM95 ; IWY03 ; LMC05 ; LP07 , blind quantum computation DKL12 , and the demonstration of quantum advantage CKDK21 . All techniques except for that in Ref. DKL12 were developed for the transmission of classical bits. For example, random single-qubit states chosen from $\{|0\rangle,|1\rangle,|\pm\rangle\}$ with |±$|\pm\rangle\equiv(|0\rangle+|1\rangle)/\sqrt{2}$ are transmitted to share a secret bit string between two parties in the original protocol BB84 of QKD. This quantum communication is replaced with the communication of coherent light in Refs. B92 ; HIGM95 ; IWY03 ; LMC05 ; LP07 . However, the transmission of classical bits is insufficient for existing verification protocols. This is why we use the technique in Ref. DKL12 . Moreover, since its proof-of-principle experiment was already demonstrated over a distance of 100 km fiber JWHXSZZLYWLXZP19 , its adoption should be preferable from a practical point of view.
@@
The remaining issue to be resolved is that the technique in Ref. DKL12 seems to be incompatible with existing verification protocols. The technique replaces the transmission of single-qubit states in a single plane of the Bloch sphere with that of coherent light, and it works when the single-qubit states are chosen uniformly at random. Although the single-qubit states used in the protocol of Ref. M20 are in the $x$-$z$ plane of the Bloch sphere, these qubits have the same random basis, i.e., the basis is not random among them. To fill in this gap, we further modify the protocol in Ref. M20 (for detail, see Protocol 1).
@@ II Verification protocol with quantum communication
As the first step to construct our physically-classical verification protocol, we propose a verification protocol with communication of qubits by modifying the protocol in Ref. M20 . The purpose of our protocol is to verify quantum computation solving any problem in ${\sf BQP}$:
@@
A promise problem $L=(L_{\rm yes},L_{\rm no})$ is in ${\sf BQP}$ if and only if there exists a uniform family $\{U_{x}\}_{x}$ of polynomial-size quantum circuits such that when $x\in L_{\rm yes}$, $\langle 0^{n}|U_{x}^{\dagger}(|1\rangle\langle 1|\otimes I^{\otimes n-1})U_{x}|0^{n}\rangle\geq 2/3$, and when $x\in L_{\rm no}$, $\langle 0^{n}|U_{x}^{\dagger}(|1\rangle\langle 1|\otimes I^{\otimes n-1})U_{x}|0^{n}\rangle\leq 1/3$. Here, $I\equiv|0\rangle\langle 0|+|1\rangle\langle 1|$ is the two-dimensional identity operator, $n$ is a polynomial in $|x|$, and $|x|$ is the length of the instance $x$.
@@
where for all $i$, $j$, and $x$, $p_{ij}^{(x)}\geq 0$, $\sum_{i<j}p_{ij}^{(x)}=1$, $c_{ij}^{(x)}\in\{1,-1\}$, and $X_{i}$ and $Y_{i}$ represent the Pauli-$X$ and $Y$ operators applied to the $i$-th qubit, respectively. Note that $N$ is a polynomial in $|x|$. Let us assume that the prover declares $x\in L_{\rm yes}$, i.e., the quantum computer outputs that the correct answer is YES. We also define certain non-negative values $a$ and $b$ such that $1\geq b-a\geq 1/f(|x|)$ for a polynomial function $f$. From the ${\sf BQP}$-hardness (more precisely, ${\sf QMA}$-completeness) of the $2$-local Hamiltonian problem KKR06 ; BL08 ; CM16 , the verifier can efficiently find $\{p_{ij}^{(x)},c_{ij}^{(x)}\}_{1\leq i<j\leq N}$ such that (i) when the prover is honest (i.e., the correct answer is indeed YES), there exists an efficiently preparable quantum state $|\eta\rangle$ whose energy $\langle\eta|H_{x}|\eta\rangle$ is at most $a$, and (ii) when the prover is malicious (i.e., the correct answer is NO), the ground-state energy is at least $b$. Since ${\sf BQP}$ is closed under complement, even when the prover declares $x\in L_{\rm no}$, the verifier can efficiently find $\{p_{ij}^{(x)},c_{ij}^{(x)}\}_{1\leq i<j\leq N}$ having the same property. The verifier in the protocol of Ref. M20 decides whether the prover is honest or malicious by measuring the energy of $H_{x}$ with the aid of a trusted third party.
@@
    Let $s^{\prime}_{k}\equiv s_{k}\oplus z_{k}\oplus h_{k}w_{k}$ for all $1\leq k\leq N$, where $z_{k}$ and $w_{k}$ are the $k$-th bits of $z$ and $w$, respectively. If $(-1)^{s^{\prime}_{i}+s^{\prime}_{j}}=-c_{ij}^{(x)}$, the verifier accepts the prover. Otherwise, the verifier rejects the prover.
There exist two differences between the original protocol in Ref. M20 and Protocol 1. First, a trusted third party is merged with the verifier in step 1. Second, the bases $\{h_{i}\}_{i=1}^{N}$ are randomly chosen for each qubit in Protocol 1, while the basis of all qubits is determined by a single random bit $h\in\{0,1\}$ in the original protocol. These differences are essential to devise our physically-classical verification protocol in the next section.
@@ III Verification of quantum computation with coherent light
    The verifier and prover conduct the following steps $N$ times. Note that when the prover is malicious, the prover can apply any completely-positive trace-preserving (CPTP) map in each step.
@@
      Let $m(\geq 8)$ be a natural number specified later. For the $j$-th repetition $(1\leq j\leq N)$, the verifier chooses a tuple $(\theta_{1}^{(j)},\ldots,\theta_{m}^{(j)})\in\{0,\pi/2,\pi,3\pi/2\}^{m}$ uniformly at random. Then the verifier sends the $m$ phase-randomized coherent states
@@
      The prover performs a quantum nondemolition (QND) measurement of the photon number on each of $m$ coherent states and obtains the measurement outcomes $\{n_{k}^{(j)}\}_{k=1}^{m}$. If $n_{k}^{(j)}\geq 1$, the prover keeps $|1_{\theta_{k}^{(j)}}\rangle$ at hand and discards the other $(n_{k}^{(j)}-1)$ photons. Let $m_{0}^{(j)}$ be the number of $k$’s such that $n_{k}^{(j)}=0$. At the end of this step, the honest prover possesses exactly $(m-m_{0}^{(j)})$ photons FNM1 .
@@
      The prover performs the interlaced 1D cluster computation (I1DC) protocol on the $(m-m_{0}^{(j)})$ photons as follows: for $l^{\prime}=1$ to $m-m_{0}^{(j)}-1$
@@
        Apply $CZ(H\otimes I)$ to the $l^{\prime}$-th and $(l^{\prime}+1)$-th photons, where $CZ\equiv I^{\otimes 2}-2|11\rangle\langle 11|$ is the controlled-$Z$ gate.
@@
        Measure the $l^{\prime}$-th photon in the Pauli-$X$ basis, and obtain the measurement outcome $o_{l^{\prime}}^{(j)}$.
@@
      The prover sends the measurement outcomes $\{o_{l^{\prime}}^{(j)}\}_{l^{\prime}=1}^{m-m_{0}^{(j)}-1}$ to the verifier. The prover keeps the unmeasured $(m-m_{0}^{(j)})$-th photon at hand and will use it as the $j$-th qubit $|\psi_{V}^{(j)}\rangle$ in step 2 of Protocol 1.
Figure 2: Schematic of our verification protocol with an honest prover. Classical communications and operations are represented by bold arrows. Non-classical communication and operations are represented by thin arrows. Each purple rectangular enclosure represents the Bell measurement. The verifier first sends phase-randomized coherent states with randomized polarization angles $\{\theta_{k}^{(j)}\}_{1\leq k\leq m,1\leq j\leq N}\in\{0,\pi/2,\pi,3\pi/2\}^{mN}$ to the prover. Then the prover performs QND measurements on all the received coherent states and sends the measurement outcomes $\{n_{k}^{(j)}\}_{k,j}$ to the verifier. The prover also discards unnecessary photons and performs the I1DC protocol by using the remaining photons (see steps (b) and (e) in Protocol 2). As a result, the prover obtains an $N$-qubit state $\otimes_{j=1}^{N}|\psi_{V}^{(j)}\rangle$, which will be used as $|\psi_{V}\rangle$ in step 1 of Protocol 1. The prover performs the quantum teleportation of the low-energy state $|\eta\rangle$ by measuring $N$ pairs of each qubits of $|\psi_{V}\rangle$ and $|\eta\rangle$ in the Bell bases and sending $\{w,z\}_{w,z\in\{0,1\}^{N}}$ to the verifier. On the other hand, the verifier checks if Eq. (5) holds for all $1\leq j\leq N$. If it does not hold, the verifier rejects the prover. If it holds, the prover chooses the pair $(i,j)$ with probability $p_{ij}^{(x)}$ and then checks whether $h_{i}=h_{j}$ (see step 3 in Protocol 1). If it is not satisfied, the prover is automatically accepted. Otherwise, the verifier calculates $(-1)^{s^{\prime}_{i}+s^{\prime}_{j}}$ by using $w$ and $z$ sent from the prover (see step 4 in Protocol 1). When it is equal to $-c_{ij}^{(x)}$, the verifier accepts the prover. On the other hand, it is not the case, the prover is rejected.
By replacing step 1 in Protocol 1 with Protocol 2, we can devise our physically-classical verification protocol (see Fig. 2). To evaluate its performance, we first derive its lower bound on the acceptance probability $p_{\rm acc}$ for the honest prover as a function of $\alpha$ and $m$. In step (b) of Protocol 2, the honest prover measures
@@
For any $1\leq l\leq m-m_{0}^{(j)}$, let $\sigma_{l}^{(j)}\in\{0,\pi/2,\pi,3\pi/2\}$ be the polarization angle of the $l$-th remaining photon at the end of step (b) in Protocol 2, i.e., the state of the $l$-th remaining photon is $(|0\rangle+e^{i\sigma_{l}^{(j)}}|1\rangle)/\sqrt{2}$ FNM3 . As shown in Ref. DKL12 , when the input states are $\{(|0\rangle+e^{i\sigma_{l}^{(j)}}|1\rangle)/\sqrt{2}\}_{l=1}^{m-m_{0}^{(j)}}$, the I1DC protocol outputs the measurement outcomes $\{o_{l^{\prime}}^{(j)}\}_{l^{\prime}=1}^{m-m_{0}^{(j)}-1}\in\{0,1\}^{m-m_{0}^{(j)}-1}$ and the single-qubit state $(|0\rangle+e^{i\varphi^{(j)}}|1\rangle)/\sqrt{2}$, where
@@
In conclusion, if Eq. (5) is satisfied for all $1\leq j\leq N$, the verifier accepts the prover with the same probability as that of Protocol 1, i.e., with probability at least $21-a/$ 2. Thus, from Eq. (8), the lower bound on $p_{\rm acc}$ of our physically-classical verification protocol is
@@
We next show the information-theoretical soundness of our physically-classical verification protocol. In other words, we give an upper bound on $p_{\rm acc}$ in the case of the malicious prover. To this end, we observe that the argument in Ref. DKL12 can be applied even in our situation. As an important point, how many photons are transmitted to the prover is randomly decided following the Poisson distribution, and hence the malicious prover cannot decide it even though any quantum operation is allowed for the malicious prover. We use this randomness to detect the malicious prover’s deviation.
Let $m_{0}^{(j)}$ and $m_{1}^{(j)}$ be the actual numbers of $k$’s such that $n_{k}^{(j)}=0$ and $1$, respectively. Although the malicious prover may not perform QND measurements in step (b) of Protocol 2, the actual photon number $n_{k}^{(j)}$ is properly defined because the phase-randomized coherent state in Eq. (6) is diagonalized in the Fock basis. Despite the malicious prover can perform any quantum operation, there exist only two cases where
@@
    for at least a single $j$, the number $\tilde{m}_{0}^{(j)}$ of vacuum states calculated from the measurement outcomes sent by the malicious prover in step (c) of Protocol 2 is at least $m_{0}^{(j)}+m_{1}^{(j)}$, and

### v2

@@ Abstract
Verification of quantum computation is a task to efficiently check whether an output given from a quantum computer is correct. Existing verification protocols conducted between a quantum computer to be verified and a verifier necessitate quantum communication to unconditionally detect any malicious behavior of the quantum computer solving any promise problem in ${\sf BQP}$. In this paper, we remove the necessity of the communication of qubits by proposing a “physically classical” verification protocol in which the verifier just sends coherent light to the quantum computer.
@@ I Introduction
Quantum computers are expected to outperform classical computers in a variety of applications, from cryptanalysis BHT97 ; S97 ; HSTX20 to physics K95 ; NPdB21 ; CDBBT24 and chemistry KJLMA08 ; GL22 ; CFGHLMW23 . The flip side of their advantages is their susceptibility to noises. Therefore, to obtain benefits from quantum computers, it is necessary to devise an efficient protocol for checking whether a quantum computer outputs a correct answer, a task called verification of quantum computation GKK19 ; EHWRMPCK20 ; KR21 . One may think that verification protocols would become useless if a sufficient number of qubits for quantum error correction G20 can be created; however, this is not the case, because it will still be necessary to check whether an implemented quantum error correction scheme faithfully works. Multiple small-scale experiments BFKW13 ; ZZCPXYYYHXCLG20 ; JWQCCLXSZM20 ; FMMD21 have already demonstrated progress toward the realization of verifiable quantum information processing.
@@
Verification protocols are evaluated in terms of five properties: (i) whether the soundness is information theoretic or computational; (ii) the type of communication required for a verifier; (iii) the number of necessary non-communicating provers, which are quantum computers to be verified; (iv) the presence or absence of a trusted third party who assuredly follows procedures in the protocols; and (v) the problems to which the protocols can be applied. Here, soundness can be information theoretic or computational. Information-theoretic soundness means that any malicious computationally unbounded prover who outputs incorrect answers can be rejected. On the other hand, computational soundness is a property for rejecting malicious quantum polynomial-time provers. Hence, the former property is stronger than the latter. As for (iii), it would, in general, be hard to guarantee that multiple provers do not communicate with each other. Therefore, the ultimate goal is to devise a protocol such that (i) the soundness is information theoretic, i.e., even if the quantum computer outputs an incorrect answer by taking superpolynomial time, it can be properly detected, (ii) the classical communication is sufficient, (iii) & (iv) a single prover is sufficient, and (v) the protocol is applicable to any problem in ${\sf BQP}$, which is a set of promise problems (i.e., problems that can be answered by YES or NO) solvable in quantum polynomial time. However, it is hard to construct such an outstanding protocol with currently known theoretical techniques. Although its impossibility has not been shown, which immediately implies ${\sf P}\neq{\sf PSPACE}$, some existing results FHM18 ; MT20 have revealed the difficulty of its construction.
@@
This situation has led to the development of numerous verification protocols M12 ; RUV13 ; M14 ; GKW15 ; HM15 ; MNS16 ; M16 ; HKSE17 ; KW17 ; FK17 ; FHM18 ; M18 ; DOF18 ; TM18 ; B18 ; FKD18 ; LMNT18 ; MTN18 ; TMMMF19 ; CGJV19 ; G19 ; GV19 ; FKD19 ; LDTAF19 ; ZH19 ; M20 ; ACGH20 ; CCY20 ; LMKO21 ; CLLW22 ; KKLMO22 ; M22 ; Z22 ; GMP23 ; LZH23 , each of which has its own advantages and disadvantages. These protocols fall into five types of approaches FN1 as summarized in Fig. 1. In terms of practicality, we stick to achieving the information-theoretic soundness and keeping the total number of provers and a trusted third party one. This is because it would be unclear how to guarantee that the prover’s computation is completed in polynomial time, that the multiple provers do not communicate with each other, and that a third party does not cooperate with a prover. In this sense, from Fig. 1, quantum communication is necessary to verify any problem in ${\sf BQP}$ with existing practical verification protocols.
@@
Figure 1: Summary of existing verification protocols. The third column indicates the type of communication required for the verifier. In the fifth column, the cross and circle marks mean that a trusted third party is unnecessary and necessary, respectively. In the sixth column, BQP means that the protocols can be applied to any problem in ${\sf BQP}$. The disadvantages of the verification protocols are highlighted in bold red. Our contribution is emphasized in bold blue.
@@
In this paper, to remove the necessity of qubit communication, we propose a “physically classical” verification protocol in which the verifier just sends coherent light to the quantum computer to be verified. The transmission of coherent light is the same as classical communication in the sense that classical bits are sent by using light in the real world, while they are absolutely different from the viewpoint of information theory. To obtain our protocol, we first modify the verification protocol in Ref. M20 such that a trusted third party is combined with the verifier, and hence qubit communication becomes necessary for the verifier. Then, to remove the qubit communication, we combine it with the technique of using coherent light as a substitute for qubits. This technique has been developed in several quantum information processing tasks such as quantum key distribution (QKD) B92 ; HIGM95 ; IWY03 ; LMC05 ; LP07 , blind quantum computation DKL12 , and the demonstration of quantum advantage CKDK21 . All the approaches except for that in Ref. DKL12 were developed for the transmission of classical bits. For example, random single-qubit states chosen from $\{|0\rangle,|1\rangle,|\pm\rangle\}$ with |±$|\pm\rangle\equiv(|0\rangle+|1\rangle)/\sqrt{2}$ are transmitted to share a secret bit string between two parties in the original QKD protocol BB84 . This quantum communication is replaced with coherent light communication in Refs. B92 ; HIGM95 ; IWY03 ; LMC05 ; LP07 . However, the transmission of classical bits is insufficient for existing verification protocols, which is why we use the technique in Ref. DKL12 . Moreover, since its proof of principle experiment has already been demonstrated over a distance of 100 km fiber JWHXSZZLYWLXZP19 , its use is preferable from a practical point of view.
@@
The remaining issue to be resolved is that the technique in Ref. DKL12 seems to be incompatible with existing verification protocols. The technique replaces the transmission of single-qubit states in a single plane of the Bloch sphere with that of coherent light, and it works when the single-qubit states are chosen uniformly at random. Although the single-qubit states used in the protocol of Ref. M20 are in the $x$-$z$ plane of the Bloch sphere, these qubits have the same random basis, i.e., the basis is not random among them. To fill in this gap, we further modify the protocol in Ref. M20 (for details, see Protocol 1).
@@ II Verification protocol with quantum communication
As the first step to constructing our physically classical verification protocol, we propose a verification protocol with communication of qubits by modifying the protocol in Ref. M20 . The purpose of our protocol is to verify quantum computation solving any problem in ${\sf BQP}$:
@@
A promise problem $L=(L_{\rm yes},L_{\rm no})$ is in ${\sf BQP}$ if and only if there exists a uniform family $\{U_{x}\}_{x}$ of polynomial-size quantum circuits such that when $x\in L_{\rm yes}$, $\langle 0^{n}|U_{x}^{\dagger}(|1\rangle\langle 1|\otimes I^{\otimes n-1})U_{x}|0^{n}\rangle\geq 2/3$ and when $x\in L_{\rm no}$, $\langle 0^{n}|U_{x}^{\dagger}(|1\rangle\langle 1|\otimes I^{\otimes n-1})U_{x}|0^{n}\rangle\leq 1/3$. Here, $I\equiv|0\rangle\langle 0|+|1\rangle\langle 1|$ is the two-dimensional identity operator, $n$ is a polynomial in $|x|$, and $|x|$ is the length of the instance $x$.
@@
where for all $i$, $j$, and $x$, $p_{ij}^{(x)}\geq 0$, $\sum_{i<j}p_{ij}^{(x)}=1$, $c_{ij}^{(x)}\in\{1,-1\}$, and $X_{i}$ and $Y_{i}$ represent the Pauli-$X$ and $Y$ operators applied to the $i$th qubit, respectively. Note that $N$ is a polynomial in $|x|$. Let us assume that the prover declares $x\in L_{\rm yes}$, i.e., the quantum computer outputs YES as the correct answer. We also define certain non-negative values $a$ and $b$ such that $1\geq b-a\geq 1/f(|x|)$ for a polynomial function $f$. From the ${\sf BQP}$-hardness (more precisely, ${\sf QMA}$-completeness) of the $2$-local Hamiltonian problem KKR06 ; BL08 ; CM16 , the verifier can efficiently find $\{p_{ij}^{(x)},c_{ij}^{(x)}\}_{1\leq i<j\leq N}$ such that (i) when the prover is honest (i.e., the correct answer is indeed YES), there exists an efficiently preparable quantum state $|\eta\rangle$ whose energy $\langle\eta|H_{x}|\eta\rangle$ is at most $a$, and (ii) when the prover is malicious (i.e., the correct answer is NO), the ground-state energy is at least $b$. Since ${\sf BQP}$ is closed under complement, even when the prover declares $x\in L_{\rm no}$, the verifier can efficiently find $\{p_{ij}^{(x)},c_{ij}^{(x)}\}_{1\leq i<j\leq N}$ having the same property. The verifier in the protocol of Ref. M20 decides whether the prover is honest or malicious by measuring the energy of $H_{x}$ with the aid of a trusted third party.
@@
    Let $s^{\prime}_{k}\equiv s_{k}\oplus z_{k}\oplus h_{k}w_{k}$ for all $1\leq k\leq N$, where $z_{k}$ and $w_{k}$ are the $k$th bits of $z$ and $w$, respectively. If $(-1)^{s^{\prime}_{i}+s^{\prime}_{j}}=-c_{ij}^{(x)}$, the verifier accepts the prover. Otherwise, the verifier rejects the prover.
There exist two differences between the original protocol in Ref. M20 and Protocol 1. First, a trusted third party is merged with the verifier in step 1. Second, the bases $\{h_{i}\}_{i=1}^{N}$ are randomly chosen for each qubit in Protocol 1, while the basis of all qubits is determined by a single random bit $h\in\{0,1\}$ in the original protocol. These differences are essential for devising our physically classical verification protocol, which is described in the next section.
@@ III Verification of quantum computation with coherent light
Figure 2: Schematic of our verification protocol with an honest prover. Classical communication and operations are represented by bold arrows. Non-classical communication and operations are represented by thin arrows. Each purple rectangular enclosure represents the Bell measurement. The verifier first sends phase-randomized coherent states with randomized polarization angles $\{\theta_{k}^{(j)}\}_{1\leq k\leq m,1\leq j\leq N}\in\{0,\pi/2,\pi,3\pi/2\}^{mN}$ to the prover. Then the prover performs QND measurements on all the received coherent states and sends the measurement outcomes $\{n_{k}^{(j)}\}_{k,j}$ to the verifier. The prover also discards unnecessary photons and performs the I1DC protocol by using the remaining photons (see steps (b) and (e) in Protocol 2). As a result, the prover obtains an $N$-qubit state $\otimes_{j=1}^{N}|\psi_{V}^{(j)}\rangle$, which will be used as $|\psi_{V}\rangle$ in step 1 of Protocol 1. The prover performs the quantum teleportation of the low-energy state $|\eta\rangle$ by measuring $N$ pairs of qubits of $|\psi_{V}\rangle$ and $|\eta\rangle$ in the Bell bases and sending $\{w,z\}_{w,z\in\{0,1\}^{N}}$ to the verifier. On the other hand, the verifier checks whether Eq. (5) holds for all $1\leq j\leq N$. If it does not hold, the verifier rejects the prover. If it holds, the prover chooses the pair $(i,j)$ with probability $p_{ij}^{(x)}$ and then checks whether $h_{i}=h_{j}$ (see step 3 in Protocol 1). If it is not satisfied, the prover is automatically accepted. Otherwise, the verifier calculates $(-1)^{s^{\prime}_{i}+s^{\prime}_{j}}$ by using $w$ and $z$ sent from the prover (see step 4 in Protocol 1). When it is equal to $-c_{ij}^{(x)}$, the verifier accepts the prover. On the other hand, if it is not the case, the prover is rejected.
@@
    The verifier and prover conduct the following steps $N$ times. Note that when the prover is malicious, the prover can apply any completely positive trace-preserving (CPTP) map in each step.
@@
      Let $m(\geq 8)$ be a natural number specified later. For the $j$th repetition $(1\leq j\leq N)$, the verifier chooses a tuple $(\theta_{1}^{(j)},\ldots,\theta_{m}^{(j)})\in\{0,\pi/2,\pi,3\pi/2\}^{m}$ uniformly at random. Then the verifier sends the $m$ phase-randomized coherent states
@@
      The prover performs a quantum nondemolition (QND) measurement of the photon number on each of the $m$ coherent states and obtains the measurement outcomes $\{n_{k}^{(j)}\}_{k=1}^{m}$. If $n_{k}^{(j)}\geq 1$, the prover keeps $|1_{\theta_{k}^{(j)}}\rangle$ at hand and discards the other $(n_{k}^{(j)}-1)$ photons. Let $m_{0}^{(j)}$ be the number of $k$’s such that $n_{k}^{(j)}=0$. At the end of this step, the honest prover possesses exactly $(m-m_{0}^{(j)})$ photons FNM1 .
@@
      The prover performs the interlaced 1D cluster computation (I1DC) protocol on the $(m-m_{0}^{(j)})$ photons as follows: for $l^{\prime}=1$ to $m-m_{0}^{(j)}-1$,
@@
        Apply $CZ(H\otimes I)$ to the $l^{\prime}$th and $(l^{\prime}+1)$th photons, where $CZ\equiv I^{\otimes 2}-2|11\rangle\langle 11|$ is the controlled-$Z$ gate.
@@
        Measure the $l^{\prime}$th photon in the Pauli-$X$ basis, and obtain the measurement outcome $o_{l^{\prime}}^{(j)}$.
@@
      The prover sends the measurement outcomes $\{o_{l^{\prime}}^{(j)}\}_{l^{\prime}=1}^{m-m_{0}^{(j)}-1}$ to the verifier. The prover keeps the unmeasured $(m-m_{0}^{(j)})$th photon at hand and will use it as the $j$th qubit $|\psi_{V}^{(j)}\rangle$ in step 2 of Protocol 1.
By replacing step 1 in Protocol 1 with Protocol 2, we can devise our physically classical verification protocol (see Fig. 2). To evaluate its performance, we first derive its lower bound on the acceptance probability $p_{\rm acc}$ for the honest prover as a function of $\alpha$ and $m$. In step (b) of Protocol 2, the honest prover measures
@@
For any $1\leq l\leq m-m_{0}^{(j)}$, let $\sigma_{l}^{(j)}\in\{0,\pi/2,\pi,3\pi/2\}$ be the polarization angle of the $l$th remaining photon at the end of step (b) in Protocol 2, i.e., the state of the $l$th remaining photon is $(|0\rangle+e^{i\sigma_{l}^{(j)}}|1\rangle)/\sqrt{2}$ FNM3 . As shown in Ref. DKL12 , when the input states are $\{(|0\rangle+e^{i\sigma_{l}^{(j)}}|1\rangle)/\sqrt{2}\}_{l=1}^{m-m_{0}^{(j)}}$, the I1DC protocol outputs the measurement outcomes $\{o_{l^{\prime}}^{(j)}\}_{l^{\prime}=1}^{m-m_{0}^{(j)}-1}\in\{0,1\}^{m-m_{0}^{(j)}-1}$ and the single-qubit state $(|0\rangle+e^{i\varphi^{(j)}}|1\rangle)/\sqrt{2}$, where
@@
In conclusion, if Eq. (5) is satisfied for all $1\leq j\leq N$, the verifier accepts the prover with the same probability as that of Protocol 1, i.e., with probability of at least $21-a/$ 2. Thus, from Eq. (8), the lower bound on $p_{\rm acc}$ of our physically classical verification protocol is
@@
We next show the information-theoretic soundness of our physically classical verification protocol. In other words, we give an upper bound on $p_{\rm acc}$ in the case of the malicious prover. To this end, we observe that the argument in Ref. DKL12 can be applied even in our situation. As an important point, how many photons are transmitted to the prover is randomly decided following the Poisson distribution, and hence the malicious prover cannot decide it even though any quantum operation is allowed for the malicious prover. We use this randomness to detect the malicious prover’s deviation.
Let $m_{0}^{(j)}$ and $m_{1}^{(j)}$ be the actual numbers of $k$’s such that $n_{k}^{(j)}=0$ and $1$, respectively. Although the malicious prover may not perform QND measurements in step (b) of Protocol 2, the actual photon number $n_{k}^{(j)}$ is properly defined because the phase-randomized coherent state in Eq. (6) is diagonalized in the Fock basis. Despite the malicious prover’s ability to perform any quantum operation, there exist only two cases where
@@
    for at least a single $j$, the actual number $m_{0}^{(j)}+m_{1}^{(j)}$ of vacuum and single-photon states is less than or equal to the threshold value $me^{-\alpha^{2}}(1+\alpha^{2}/2)$ in step (d) of Protocol 2, and